OpenStudy (anonymous):

Find the non-extraneous solutions of the square root of the quantity of x plus 3 minus 4 equals x minus 1. x = −3 and x = −2 x = −2 x = −3 x = 5

OpenStudy (driftracer305):

i got x = 2 or 1......

OpenStudy (aum):

Not at all clear what the equation is. Can you use the equation editor to retype? Or post a screenshot of the question.

OpenStudy (driftracer305):

he said it was sqrt of x+3 -4 that equals to x-1.. solve for x @aum

OpenStudy (aum):

It could be sqrt(x+3) - 4 = x - 1 also. Not clear where the square root stops.

OpenStudy (driftracer305):

\[\sqrt{x+3-4} = x-1\] \[\sqrt{x-1} = x-1\] \[x-1 = x^2-2x+1\] \[(x-2)(x-1)\] \[x=2 or 1....... both work\]

OpenStudy (aum):

I think the equation is \(\Large \sqrt{x+3} - 4 = x -1\).

OpenStudy (aum):

The first solution works for the equation in my previous reply.

OpenStudy (driftracer305):

so it is obvious @aum ........ i think your equation fits the choices here..... good logic..... i was able to solve it.....

OpenStudy (anonymous):

so A?

OpenStudy (driftracer305):

@thedukelife ... it is A

OpenStudy (anonymous):

thanks

OpenStudy (driftracer305):

np... glad to help

OpenStudy (anonymous):

Which of the following represents 6 x to the 3 fourths power in radical form? cubed root of 6 x to the fourth power fourth root of 6 x cubed 6 cubed root of x to the fourth power 6 fourth root of cubed

OpenStudy (anonymous):

what about this..?

OpenStudy (driftracer305):

k first of all start this as a new question pls....... i get confused

OpenStudy (anonymous):

ok